A company which produces batteries claims that the life expectancy of their batteries is 90 hours. In order to test the claim, a consumer interest group tested a random sample of 40 batteries. The test resulted to a mean lif expectancy of 87 hours. Usoing a 0.05 level of significance, can it concluded that the life expectancy of their batteries is less than 90 hours? Assume that the population standard deviation is known to be 10 hours.


1
Expert's answer
2021-05-11T14:13:37-0400

"H_0:" The mean life expectancy of their batteries is 90 hours, i.e. "H_0: \u03bc= 90"


"H_1:" The mean life expectancy of their batteries is less than 90 hours, i.e. "H_1: \u03bc < 90"


n=40 (large sample)

"\bar{x}=87 hours"


σ=10 hours


μ= 90


Level of significance"= \u03b1 =0.05"


Since the sample size is large we can use the test statistic as Z


Test statistics: 


"Z = \dfrac{\bar{x}-\mu}{\frac{\sigma}{\sqrt{n}}}"



            "= \dfrac{87-90}{ \frac{10}{\sqrt{ 40}}}"



           "= \dfrac{-3}{ 1.58}"


           "= - 1.90"


P-value"= P(Z < - 1.90)= P(Z > 1.90)"


                   "=0.5 \u2013 0.4713" {Value taken from standard normal table}


                   =0.0287


   0.0287 < 0.05


   P-value < level of significance


  Reject "H_0" .


Therefore, we have sufficient evidence to conclude that the mean life expectancy of their batteries is less than 90 hours .